# S2-Diffusion: Generalizing from Instance-level to Category-level Skills in Robot Manipulation

### v2

@@ Abstract
Recent advances in skill learning has propelled robot manipulation to new heights by enabling it to learn complex manipulation tasks from a practical number of demonstrations. However, these skills are often limited to the particular action, object, and environment instances that are shown in the training data, and have trouble transferring to other instances of the same category. In this work we present an open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) which enables generalization from instance-level training data to category-level, enabling skills to be transferable between instances of the same category. We show that functional aspects of skills can be captured via a promptable semantic module combined with a spatial representation. We further propose leveraging depth estimation networks to allow the use of only a single RGB camera. Our approach is evaluated and compared on a diverse number of robot manipulation tasks, both in simulation and in the real world. Our results show that S2-Diffusion is invariant to changes in category-irrelevant factors as well as enables satisfying performance on other instances within the same category, even if it was not trained on that specific instance. Full videos of all real-world experiments are available in the supplementary material.
@@ I INTRODUCTION
Figure 1: Our open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) not only efficiently completes the task at hand but also enables the generalization of the same skill across diverse contexts and task variations. For instance, a rice-scooping policy can be directly transferred to scoop cereal without any fine-tuning.
@@
Large pre-trained Visual-Language-Action (VLA) models [4, 5, 6] generalize simple skills such as pick-and-place over a wide range of environments and objects. However, more complex non-prehensile manipulation such as scooping, mug-flipping, cooking shrimp or opening a bottle with a bottle opener stay elusive for such large general models. Recent imitation learning approaches address learning of instances of challenging manipulation tasks  [1, 2, 7, 8, 9, 10, 11] but the integration with semantic knowledge in highly-complex manipulation tasks remains a challenge [9]. The aforementioned methods often rely on raw perceptual features and environmental conditions, limiting their applicability to the instances observed during the training. Training a skill via imitation learning, such as a diffusion policy [9], relies on expert demonstrations that often do not cover several instances of the same task. We show that when trained only on a particular instance of a task - such as wiping red scribbles from a whiteboard - the skill fails to transfer already when the scribbles are now green, even if the required action and environment for wiping is exactly the same. This is because the policy did not learn a whiteboard-wiping category skill but a single instance of this category namely red-whiteboard-wiping.
@@
Motivated by the above challenges, we present a novel approach that integrates spatial prediction with semantic segmentation features from large pre-trained models [12, 13] to generalize from expert demonstrations on a single instance of a task to its category-level skill – such as wiping different color scribbles from the whiteboard or scooping different granular material from bowl to bowl. Our method uses a single RGB camera view combined with the proprioceptive information of the robot. As shown in Fig. 1, we extract semantic information using the prompted foundation model [14] and combine it with a depth estimation foundation model [15] to obtain Spatial-Semantic observations for the visuomotor diffusion policy learning framework. This allows for invariance to task-irrelevant factors such as background and object textures as well as the capability to generalize from instance-level to category-level skills. Our key contributions are threefold:
  1. 1.
    We propose to endow visuomotor diffusion policies with spatial-semantic understanding to enable generalization from instance-level to category-level skills in robot manipulation tasks.
  2. 2.
    We introduce an efficient representation of the spatial-semantic information via a combination of vision foundation models. The overall framework is real-time viable and requires only a single RGB camera and the robot’s proprioceptive observations.
  3. 3.
    Our extensive experiments evaluate the method on a set of robotics manipulation tasks in simulation and the real world, demonstrating the ability to learn generalizable and effective robotic manipulation policies.
All real-world experimental videos ($174$) are available in the supplementary material.
@@ II RELATED WORK
II-A Visual Feature Based Imitation Learning
In recent studies, visual imitation learning methods [16, 17, 18, 19] have shown great potential in diverse robot manipulation tasks. The recent work of Zhu et al. [3] demonstrates the potential of object-based representations for manipulation by using a conventional closed-set object detector. In [20], the authors validate that large-scale video generative pre-training is able to effectively benefit visual robot manipulation learning. However, in their works, the finetuned policy is not demonstrated to generalize to novel instances of the task. 3D vision based imitation learning [21, 22, 23, 24] has shown to improve performance but typically requires multiple posed views with RGB-D sensors to reconstruct the scene. Another work [25] also ulitizes pretrained VLMs to provide scene representation for robot tasks, but their approach needs to construct a 3D neural feature field by scanning part of the scene. In our paper, we propose a visuomotor imitation learning method that takes advantage of object-aware semantic representation to focus on task-relevant objects for training a visuomotor policy.
Recently, Wang et al. [26] introduce a novel framework that enhances generalization capabilities by incorporating explicit spatial and semantic information into policy learning. This method uses multi-view RGB-D observations to generate 3D descriptor fields and then constructs semantic fields that highlight task-relevant parts of objects (e.g., a knife handle or soda can tab). By combining 3D geometric and semantic information, this approach achieves significant improvements in generalization at the cost of requiring multiple RGB-D cameras.
II-B Vision Language Models in Robotics
Vision-Language Models (VLMs), trained on vast amounts of internet data, have found widespread use in robotic applications [27, 28, 29]. Some works [30, 31, 32, 26] have explored leveraging 2D foundation vision models to construct open-vocabulary 3D representations. These approaches involve utilizing visual features from pretrained foundation models like CLIP [27] or SAM [33] to scene representations that benefit learning downstream robot policy. The most recent works [12, 34, 35] demonstrate superior open-vocabulary classification ability which succeed in generalizing to unseen classes. Pretrained foundation models are also utilized by either being prompted or fine-tuned to generate high-level task plans [36, 31] and to shape reward functions [37, 38, 39] for reinforcement learning. In [40], the authors introduce a multimodal diffusion transformer policy for solving long-horizon tasks. In contrast, our approach concentrates on improving generalization ability for open-vocabulary manipulation policy. MOKA [41] addresses robotic manipulation by using VLMs to mark open-world keypoint affordances. The core idea behind both MOKA and our method is similar, that is guiding the policy to focus on task-relevant objects. However, our method utilizes a more generalizable semantic representation and also depth information for the entire scene.
II-C Vision Language Actions in Robotics
Large policies or Vision-Language Actions (VLAs), pretrained on a combination of Internet-scale data and diverse robot demonstrations, is another promising research direction for generalizable policy learning in robotics. Recent works [42, 5, 43, 44] propose to co-finetune robot trajectory data alongside vision-language data, demonstrating impressive generalization capabilities with novel objects and instructions. However, these are complex and computationally costly attempts at general manipulation that while providing encouraging results at task reasoning have so far not been able to compete with imitation-based skill learning on non-prehensile and contact-rich manipulation tasks. To enhance generalization ability, [45] further proposes a cross-embodied robot policy trained on the diverse robot dataset to control robots with varying observation and action types. For comparison, our method focuses on policy generalization across different tasks by concentrating solely on the segmentation masks of target objects.
@@ III Towards Generalizable Robotic Skills
We start with a question: "What does it mean for a skill to be generalizable?". Commonly, "mug flipping" trained on a specific mug with distinct appearance and geometry will not generalize to flipping other mugs. Instead, the learned skill remains instance-specific, as neural network-based methods optimize the training loss without explicit interpretability of the learned features. Consequently, if these (self-selected) features are absent in another mug instance, the skill will fail to perform effectively.
One approach to mitigate this limitation is to explicitly learn features that are shared across all instances within a category. For example, a "mug" could be characterized by shared geometric features, such as a cylindrical body and a handle. Works such as ShapeNet [46] provide large-scale repositories of 3D shapes categorized into object classes, enabling methods that are able to learn category-level features that can be generalized over individual instances [47]. These approaches enable generalization across objects within the same class while preserving distinctions between categories.
This object-centric approach has proven effective for tasks involving well-categorized objects like mugs, shoes, pens, or soda cans. However, tasks like wiping or scooping present unique challenges that cannot be addressed solely by these methods. In this work, we follow a more action-centric approach, that considers the semantic functionality of the skill rather than specific object categories.
In this context, the concept of affordances [48] provides a useful lens to interpret our action-centric approach. Affordances, which describe the potential actions that an object or environment can support, emphasize functionality over rigid categorization. While our work focuses on learning lower-level skills and contact-rich tasks, which are difficult to reduce to affordances, it shares a similar philosophy by focusing on the semantic functionality of skills rather than specific object categories.
III-A Functional and Semantic Abstractions
When considering robotic skills, it is essential to move beyond instance-specific implementations and emphasize a leveled abstraction of skills. This facilitates generalization across instances and categories, aligning with the notion that skills can share functional similarities despite differing in object appearances, materials, or environmental contexts. In Fig. 2 we show an example of such a skill levels - depicting the functional goals on the top, able to be achieved with different category-level skills and performed on instances of that category.
Figure 2: Skill level example for flipping and scooping. The functional goal is stated as an abstract action on the top row, second row shows examples of category-level skills that fall under this functional similarity. The bottom row shows Instances of the respective categories.
As the example in Fig. 2 shows, the skill flipping represents a generic functional task: "the act of inverting an object". Within this abstraction, mug flipping is a category-level skill that applies to any object classified as a "mug". However, when trained specifically on a single mug (e.g., "a blue cylindrical mug with a narrow handle"), the skill becomes an instance-specific skill, which may not generalize to mugs with varying shapes, textures, or weights. Similarly, the skill scooping provides another illustrative example. At its core, scooping refers to the functional action of transferring a granular or semi-solid material from one location to another. Category-level skill: Bowl-to-bowl scooping captures the general action of transferring materials between two containers, encompassing a wide range of granular and semi-solid materials. Instance-level skills: Tasks such as rice scooping or cereal scooping are more specific, as they have different visual appearances as well as grain size and density.
In this work, we define an instance-level skill as a skill learned from specific demonstrations that are tightly coupled to a particular object, material, or environmental setup. For example, a rice-bowl-to-bowl scooping policy trained exclusively on a dataset containing only rice would constitute an instance-level skill, as the robot may fail to generalize when attempting to scoop a different granular material, such as cereal or sand. On the other hand, a category-level skill is a more abstract representation of a skill that remains valid across multiple instances within a category. For example, bowl-to-bowl scooping encompasses a variety of granular materials, such as rice, cereal, or lentils, enabling the robot to execute the task reliably across these different instances. Category-level skills incorporate a more generalized functional understanding, allowing the robot to adapt to new but related scenarios without explicit retraining.
In practice, all training data - in the real world - will come from instance-specific setups for a category-level skill. One way to achieve generalization is simple to present the imitation learning method with a wide range of instances in order to cover the whole category and achieve generalization. This approach, however, has the drawback of being very labor-intensive and therefore costly when done in the real world. We address this shortcoming by focusing on Spatial-Semantic features that can be extracted from a single instance but are still informative enough so that generalization can happen. In other words when we learn a visuomotor diffuson policy with Spatial-Semantic observations - instead of RGB - we can generalize from instance-level to its category-level skill.
@@ IV PROBLEM FORMULATION
Let $\Psi$ be the space of RGB images. Consider a manipulation task given the language instruction of the target object $\mathcal{L}$ (e.g., "coffee mug"). Our goal is to learn a generalizable imitation learning policy by leveraging spatial-semantic representation from pretrained Vision-Language models.
We consider a robotic task as a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma)$ of states, actions, transition probability, reward, initial state distribution, and discount factor. We assume access to a dataset $D=\{\tau_{i}\}_{i}^{N}$ of $N$ demonstrated expert trajectories $\tau_{i}=\{(s_{0},a_{0}),...,(s_{T_{i}},a_{T_{i}})\}$ for the task. Our method uses a denoising diffusion process [9] to learn imitation policy $\pi$ from $D$. To generalize from individual instances to unseen instances from the same category, we propose to build a spatial-semantic representation. In this paper, $\mathcal{S}$ is the state space composed of visual spatial-semantic representation $f_{v}$ and robot proprioception states $q$. $\mathcal{A}$ is the action space of robot end-effector commands. We aim to learn a policy $\pi_{\theta}(a|s):\mathcal{S}\rightarrow\mathcal{A}$ with parameters $\theta$ that predicts action $a$ according to current state $s$ for an MDP by leveraging the prior experience contained in the dataset $D$.
@@ V-A Diffusion for Robot Skill Learning
$a^{k-1}=\alpha_{k}(a_{k}-\gamma_{k}\epsilon_{\theta}(a^{k},o,k))+\sigma_{k}\mathcal{N}(0,I),\$ (1)
@@
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{0}+\epsilon^{k},o,k)),\$ (2)
@@
V-B Open-Vocabulary Spatial-Semantic Representation
We aim to design spatial-semantic-aware representation that is leveraged for the observation of the above denoising probabilistic model. We utilize two pretrained VLMs, Grounded-SAM2 [14] and DepthAnythingV2 [13], for open-vocabulary semantic segmentation and depth map estimation respectively. We use Grounded-SAM2 model to perform zero-shot semantic segmentation, leveraging CLIP-based [27] mask classification to segment unseen classes. We combine the extracted features to construct a spatial-semantic representation that is leveraged as the input for the visuomotor diffusion policy.
We utilize Grounded-SAM2 to segment an image into a set of semantic masks $(z_{1},z_{2},...,z_{n})$ from visual observations based on text descriptions. We apply pixel-wise maximum pooling for each pixel location $(i,j)$ across all segmentation masks:
$z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j),$ (3)
@@ 13                 $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
To improve the spatial reasoning of the imitation policy, we propose to incorporate depth map of the task space into its semantic representation. Specifically, the input image is also processed separately with a pretrained 335M-parameter DepthAnythingV2 model [13] that shows promising performance in fine-grained details. The Depth-Anything model predicts depth estimates relative to the input RGB observations rather than absolute values. This can result in inconsistencies and inaccuracies over extended tasks or manipulation horizons. To address this, we normalize the depth maps during both training and evaluation, ensuring consistency across diverse scenes and improving the model’s robustness and reliability.
The resulting semantic and spatial feature vectors are concatenated along the channel dimension to form the spatial-semantic representation. Upon generating the spatial-semantic representation, we leverage it as input for a visuomotor diffusion policy [9].
Figure 4: Simulated Tasks. We perform evaluations on five single-stage tasks from a large-scale simulation framework RoboCasa [52]: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda. RoboCasa provides 50 expert demonstrations for each of the manipulation tasks. In the demonstration datasets for each task, we assume variations in the color and shape of the target object, as well as differences in the background environment.
TABLE I: Success Rate of Simulation Experiments

### v3

@@ Abstract
Recent advances in skill learning has propelled robot manipulation to new heights by enabling it to learn complex manipulation tasks from a practical number of demonstrations. However, these skills are often limited to the particular action, object, and environment instances that are shown in the training data, and have trouble transferring to other instances of the same category. In this work we present an open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) which enables generalization from instance-level training data to category-level, enabling skills to be transferable between instances of the same category. We show that functional aspects of skills can be captured via a promptable semantic module combined with a spatial representation. We further propose leveraging depth estimation networks to allow the use of only a single RGB camera. Our approach is evaluated and compared on a diverse number of robot manipulation tasks, both in simulation and in the real world. Our results show that S2-Diffusion is invariant to changes in category-irrelevant factors as well as enables satisfying performance on other instances within the same category, even if it was not trained on that specific instance. Project website: https://s2-diffusion.github.io.
Index Terms:
Imitation Learning, Learning from Demonstration, Deep Learning in Grasping and Manipulation.
@@ I Introduction
Figure 1: Our Spatial-Semantic Diffusion policy (S2-Diffusion) not only efficiently completes the task at hand but also enables the generalization of the same skill across diverse contexts and task variations.
@@
Large pre-trained Visual-Language-Action (VLA) models [4, 5, 6] generalize simple skills such as pick-and-place over a wide range of environments and objects. However, more complex non-prehensile manipulation such as scooping, mug-flipping, cooking shrimp or opening a bottle with a bottle opener stay elusive for such large general models. Recent imitation learning approaches address learning of instances of challenging manipulation tasks  [2, 7, 8, 9, 10] but the integration with semantic knowledge in highly-complex manipulation tasks remains a challenge. The aforementioned methods often rely on raw perceptual features and environmental conditions, limiting their applicability to the instances observed during the training. Training a skill via imitation learning, such as a diffusion policy [2], depends on expert demonstrations that often do not cover several instances of the same task. We show that when trained only on a particular instance of a task - such as wiping red scribbles from a whiteboard - the skill fails to transfer already when the scribbles are now green, even if the required action and environment for wiping is exactly the same. This is because the policy did not learn a whiteboard-wiping category skill but a single instance of this category namely red-whiteboard-wiping.
@@
Motivated by the above challenges, we present a novel approach that integrates spatial prediction with semantic segmentation features from large pre-trained models [11, 12] to generalize from expert demonstrations on a single instance of a task to its category-level skill – such as wiping different color scribbles from the whiteboard or scooping different granular material from bowl to bowl. Our method uses a single RGB camera view combined with the proprioceptive information of the robot. As shown in Fig. 1, we extract semantic information using the prompted foundation model [13] and combine it with a depth estimation foundation model [14] to obtain Spatial-Semantic observations for the visuomotor diffusion policy learning framework. This allows for invariance to task-irrelevant factors such as background and object textures as well as the capability to generalize from instance-level to category-level skills. Our key contributions are threefold: i) We propose to endow visuomotor diffusion policies with spatial-semantic understanding to enable generalization from instance-level to category-level skills in robot manipulation tasks. ii) We introduce an efficient representation of the spatial-semantic information via a combination of vision foundation models. The overall framework is real-time viable and requires only a single RGB camera and the robot’s proprioceptive observations. iii) Our extensive experiments evaluate the method on a set of robotics manipulation tasks in simulation and the real world, demonstrating the ability to learn generalizable and effective robotic manipulation policies. All real-world experimental videos ($246$) are on the project website.
@@ II RELATED WORK
Visual Feature Based Imitation Learning. Visual imitation learning methods [15, 16] have shown strong potential in diverse robot manipulation tasks. While Zhu et al. [3] and Wu et al. [17] highlight the benefits of object-centric and large-scale video pretraining, they do not demonstrate generalization to novel task instances. 3D vision-based methods [18, 19, 20] improve generalization but require costly multi-view RGB-D setups or scene scanning. In contrast, our method uses a semantic-aware representation from a single RGB image. A concurrent work, SAM2Act [21], introduces a multi-view transformer-based robotics policy for improved visual feature representation and multitask generalization, with SAM2Act+ further incorporating a memory-based architecture for episodic recall in spatial memory-dependent manipulation tasks. Closest to our work, Wang et al. [22] integrate explicit spatial-semantic information into policy learning, but rely on multi-view 3D descriptor fields, whereas we achieve similar benefits without extra hardware.
Vision-Language Models in Robotics. Vision-Language Models (VLMs) trained on large internet corpora [23, 24] have become prominent in robotics. Recent works [25, 22] leverage 2D foundation models like CLIP [23] and SAM [26] to construct open-vocabulary 3D representations. Newer models [11, 27] improve open-world generalization. VLMs are also employed to generate high-level task plans [28, 29] or reward functions [30, 31]. Reuss et al. [32] introduce multimodal diffusion policies for long-horizon tasks. MOKA [33] uses keypoint affordances for manipulation, but we enhance generalization further by incorporating both semantic and depth information.
Vision-Language Actions in Robotics. Vision-Language Actions (VLAs) pretrained on large-scale data [34, 5, 35] show strong task reasoning but remain computationally intensive and struggle with fine-grained manipulation. Cross-embodied policies [36] attempt to scale generalization across robots. Instead, our method focuses on compact, efficient policies by leveraging open-vocabulary segmentation masks for policy generalization across manipulation tasks.
Figure 2: Skill abstraction hierarchy for flipping and scooping tasks.
@@ III Towards Generalizable Robotic Skills
What does it mean for a skill to be generalizable? Commonly, skills like "mug flipping" learned on a specific instance do not transfer to new instances with different appearance or geometry. Neural policies often overfit to instance-specific features, limiting their applicability. One solution is to explicitly learn features shared across a category. For instance, "mugs" can be characterized by a cylindrical body and handle. Datasets like ShapeNet [37] facilitate learning such category-level features.Object-centric representations work well for defined object classes (e.g., mugs, shoes), however, they struggle with tasks involving materials or actions, such as wiping or scooping. Instead, we advocate an action-centric perspective, focusing on semantic functionality over object identity.
@@
Skills should be abstracted beyond specific instances. As illustrated in Fig. 2, a functional goal like scooping encompasses various category-level skills and their specific instances. For example, scooping represents transferring granular or semi-solid materials. A category-level skill like bowl-to-bowl scooping covers materials such as rice, cereal, or lentils, whereas instance-level policies (e.g., rice scooping) often fail to generalize. In real-world robotics, demonstrations naturally reflect instance-level setups. Generalization could be attempted by collecting demonstrations across many instances, but this is costly and labor-intensive. Instead, we propose leveraging Spatial-Semantic features extracted from a single instance. By replacing raw RGB inputs with spatial-semantic observations, we enable visuomotor diffusion policies to generalize from instance-level training to category-level execution.
@@ IV PROBLEM FORMULATION
Our goal is to learn a generalizable imitation learning policy by leveraging spatial-semantic representation from pretrained Vision-Language models. We assume access to a dataset $D=\{\tau_{i}\}_{i}^{N}$ of $N$ demonstrated expert trajectories $\tau_{i}=\{(o_{0},a_{0}),...,(o_{T_{i}},a_{T_{i}})\}$ for the task. Our method uses a denoising diffusion process [2] to learn imitation policy $\pi$ from $D$. To generalize from individual instances to unseen instances from the same category, we propose to build a spatial-semantic representation. $\mathcal{O}$ is the observation space composed of visual spatial-semantic representation $f_{v}$ and robot proprioception states $q$. $\mathcal{A}$ is the action space of robot end-effector commands. We aim to learn a policy $\pi_{\theta}(a|o):\mathcal{O}\rightarrow\mathcal{A}$ with parameters $\theta$ that predicts action $a$ according to current observation $o$ by leveraging the prior experience contained in the dataset $D$.
@@ V-A Diffusion for Robot Skill Learning
$a^{k-1}=\alpha_{k}(a_{k}-\gamma_{k}\epsilon_{\theta}(a^{k},o,k))+\sigma_{k}\mathcal{N}(0,I),\\$ (1)
@@
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{0}+\epsilon^{k},o,k)),\\$ (2)
@@ 14        $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
V-B Open-Vocabulary Spatial-Semantic Representation
We aim to design spatial-semantic-aware representation that is leveraged for the observation of the above denoising probabilistic model. We utilize two pretrained VLMs, Grounded-SAM2 [13] and DepthAnythingV2 [12], for open-vocabulary semantic segmentation and depth map estimation respectively. We use Grounded-SAM2 model to perform zero-shot semantic segmentation, leveraging CLIP-based [23] mask classification to segment unseen classes. We combine the extracted features to construct a spatial-semantic representation that is leveraged as the input for the visuomotor diffusion policy. We utilize Grounded-SAM2 to segment an image into a set of semantic masks $(z_{1},z_{2},...,z_{n})$ from visual observations based on text descriptions. We apply pixel-wise maximum pooling for each pixel location $(i,j)$ across all segmentation masks:
$z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j),$ (3)
To improve the spatial reasoning of the imitation policy, we propose to incorporate depth map of the task space into its semantic representation. Specifically, the input image is also processed separately with a pretrained 335M-parameter DepthAnythingV2 model [12] that shows promising performance in fine-grained details. The model predicts depth estimates $z_{d}$ relative to the input RGB observations rather than absolute values. This can result in inconsistencies and inaccuracies over extended tasks or manipulation horizons. To address this, we normalize the depth maps during both training and evaluation, ensuring consistency across diverse scenes and improving the model’s robustness and reliability. The resulting semantic and spatial feature vectors are concatenated along the channel dimension to form the spatial-semantic representation $z=z_{f}\oplus z_{d}$, where $z\in\mathbb{R}^{(C_{s}+C_{d})\times H\times W}$ with $C_{s},C_{d}$ representing the channel dimensions of the semantic and depth features. Upon generating the spatial-semantic representation, we leverage it as input for a visuomotor diffusion policy [2].
Figure 4: Simulated Tasks. We perform evaluations on six single-stage tasks from a large-scale simulation framework RoboCasa [41]: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda, TurnOnStove, and two tasks in SAPIEN simulator: HangMug, InsertPencil.
TABLE I: Success Rate of Simulation Experiments in RoboCasa